(Directions: Let x represent one number and let y represent the other number. Use the given conditions to write a system of equations. Solve the system and find the numbers.)
The sum of two numbers is 2. If one number is subtracted from the other, the result is 8. Find the numbers.

Answer by checkley77(12844)

X+Y=2 OR X=2-Y
X-Y=8
(2-Y)-Y=8
2-2Y=8
-2Y=8-2
-2Y=6
Y=6/-2\Y=-3 ANS.
X-3=2
X=2=3
X=5 ANS.
PROOF:
5-(-3)=8
5+3=8
8=8